Write the product in standard form.
( 9 + 5i)( 9 + 8i)

Found 2 solutions by MathLover1, DrBeeee:
Answer by MathLover1(20850)
Answer by DrBeeee(684)

Given:
(1) (9+5i)*(9+8i)
You multiply this the same as you FOIL two binomials, treating the real and imaginary parts as if they were two different variables.
(2) (9+5i)*(9+8i) = 9*9 + 9*8i + 5i*9 + 5i*8i or after multiplying
(3) (9+5i)*(9+8i) = 81 + 72i + 45i + 40i^2 and after adding the imaginary parts and using the fact that i^2 = -1 we get
(4) (9+5i)*(9+8i) = 81 + 117i - 40 or after adding the real parts we get
the final answer
(5) (9+5i)*(9+8i) = 41 + 117i